OpenStudy (anonymous):

 What is the simplified form of sin(x + p)? a. cos x b. sin x c. –sin x d. –cos x

OpenStudy (anonymous):

p is what pie ??

OpenStudy (vishweshshrimali5):

yes...... sin(x+\(\pi\)) = -sin x..................

OpenStudy (anonymous):

\[\sin(\pi + x) =- \sin(x)\]

OpenStudy (vishweshshrimali5):

Cheating............ :P

Parth (parthkohli):

\[\sin(a + b) = \sin a\cos b + \cos a \sin b \]

OpenStudy (vishweshshrimali5):

Yes........ now in Parth's formula put a = x and b = \(\pi\)

OpenStudy (anonymous):

I'm confused D:

OpenStudy (vishweshshrimali5):

Why////////

OpenStudy (anonymous):

idk i just am, so is it -sinx then?

OpenStudy (vishweshshrimali5):

Yes...... but If u are having some doubt then please ask

OpenStudy (anonymous):

okay, well thank you(: